# Matching with Generalized Sequential Reserves

### v2

@@ Abstract.
This paper introduces a general theory of sequential reserves by formulating a parameterized family of generalized sequential (GS) choice rules for institutions consisting of multiple divisions (or reserve categories). We show that the cumulative offer mechanism (COM) is the unique stable and strategy-proof mechanism when institutions’ choice rules are in the GS family. Choice rules in real-life market design applications appear naturally in the GS family. We present two such real-world applications: affirmative action implementations in India and Chinese high school admissions.
@@ 1. Introduction
Many real-life allocation problems involve matching individuals with institutions that are made up of multiple divisions such as firms, hospitals, or schools. In rationing healthcare resources (Pathak, Sönmez, Ünver, and Yenmez (2024) and Aziz and Brandl (2024)) and the allocation of public goods such as seats in prestigious engineering schools (Baswana, Chakrabarti, Chandran, Kanoria, and Patange (2019), Aygün and Turhan (2022)) and government jobs (Sönmez and Yenmez (2022)) available resources are divided into different reserve categories, each with a subset of individuals eligible for that category. Divisions (or reserve categories) are endowed with their own choice rules. In some applications, divisions’ choice rules are induced from a strict ranking of alternatives, by always selecting the q-best elements whenever available. Such choice rules are called q-responsive. Similarly, in many rationing practices, there is a baseline priority ordering on applicants, and resource units are assigned according to this baseline priority order in each reserve category, also depending on the eligibility of candidates. However, in many other allocation or rationing practices, divisions’ choices are more complicated. To make it precise, consider a division that selects applicants on the basis of their merit scores but is also required to select minimum numbers of disabled individuals and women applicants, whenever possible. This choice rule is obviously not responsive.
@@
Often, divisions are interconnected in that the number of positions available in one division depends on how many applicants other divisions hire. This creates inter-divisional constraints across the institution. For example, consider a business school whose departments, such as marketing, economics, and finance, are in the job market to hire new faculty. Suppose that the marketing department hires first, followed by economics, which is then followed by finance. The number of available positions for finance might be increased if either the marketing or the economics department hires fewer faculty than they originally planned, given that the total budget of the school is not exceeded. Roth and Peranson (1999) were the first to address interdivisional constraints of this kind in matching theory. They introduced the notion of “reversion" in the context of the National Residency Matching Program (NRMP), which matches doctors with hospital programs. Reversion is a practice where unfilled positions in one hospital program can be transferred to other programs while respecting the hospital’s overall capacity limits. In their analysis, Roth and Peranson (1999) identified reversion as a significant complication in the NRMP’s matching process.
@@
There exists a significant body of literature that examines the design and implementation of soft reserve policies in rationing and resource allocation problems. The specific method of soft reservation implementation has important distributional consequences. In the context of admissions to technical universities in India, Aygün and Turhan (2022) introduced a soft reserve policy that reverts unfilled slots reserved for particular groups to the open category, which is filled before reserve categories by updating the position categories’ capacities and re-running the choice rule from scratch. In a similar framework, Aygün and Turhan (2025) developed forward transfer choice rules in which reserved slots that are not filled are allocated as open category after the reserve categories are filled. Hu and Zhang (2025) present various soft reserve policies specifically currently in use for high school admissions in China.
This paper studies the choice problem of multi-divisional institutions that face cross-divisional constraints. We introduce a family of choice rules termed generalized sequential (GS) rules, which incorporate two fundamental generalizations. First, these rules allow divisions to employ substitutable and size-monotonic choice rules rather than constraining them to simpler choice rules based on given ranking of alternatives. Second, these rules facilitate a more comprehensive implementation of soft reserves, as defined through transfer functions. Specifically, we present a parameterized family of choice rules constructed from choice rules that satisfy the substitutes property and size monotonicity—two of the most prominent properties in the market design literature. Although choice rules in the GS family may fail the substitute condition and size monotonicity, we show in our Theorem 1 that every choice rule in the GS family has a completion that satisfies the substitute condition and size monotonicity. A direct corollary of Theorem 1 is that the generalized deferred acceptance mechanism is the unique mechanism that is stable and strategy-proof in centralized markets.
@@
The choice rules in the GS family have the following structure: Each institution has a number of divisions, each of which has a certain number of available positions. There is a predetermined order of precedence over divisions according to which the institution fills the positions of its divisions. Each division is endowed with its own choice rule. A division’s choice rule has two inputs: the set of available options and its capacity. Both inputs depend on the choices made by the preceding divisions. The set of available options can be thought of as the set of remaining options from the choices of divisions that precede it. The capacity of each division is an increasing function of the number of unfilled positions of the divisions that precede it, given by an exogenously specified transfer function. The aggregate choice rule of an institution is then defined as the union of its divisions’ choice rules. The collection of divisions’ choice rules and the transfer function identify an aggregate choice rule.
@@
The choice rule for each division satisfies the substitutes and the size monotonicity properties. The substitutes condition (Kelso and Crawford (1982) and Gul and Stacchetti (1999)) and size monotonicity (Alkan and Gale (2003) and Hatfield and Milgrom (2005)) impose restrictions on the choice function when the set of available alternatives changes while the capacity is fixed. A choice rule satisfies the substitute property if there are no two alternatives $x$ and $y$ that are complementary, in the sense that gaining access to $x$ makes $y$ desirable. A choice rule is size monotonic if it chooses weakly more alternatives whenever the set of available alternatives expands.
The choice rule of each division also satisfies the quota monotonicity property, which is similar to the notion of choice rule expansion defined in Chambers and Yenmez (2017). [^1] The quota monotonicity property is a stronger condition than the expansion property of Chambers and Yenmez (2017). The quota monotonicity property imposes restrictions on the choice rule when the capacity of the division changes, while the set of available alternatives remains unchanged. A choice rule satisfies quota monotonicity if
@@
Divisions’ choice rules with initial and increased capacities can be regarded as two different choice rules that satisfy the substitutes condition and size monotonicity. The quota monotonicity property imposes a tight relationship between them. The choice rule with the increased capacity is an expansion of the choice rule with the initial capacity.
@@
We formulate cross-division constraints via transfer functions. Given the order of precedence over divisions, the capacity of a division is an increasing function of the number of vacancies in the divisions that precede it. We require that the transfer function be monotone (Westkamp (2013)). That is, it satisfies the following conditions: $(1)$ whenever there are weakly more seats unfilled in every division preceding the $j^{th}$ division, weakly more slots should be available for the $j^{th}$ division, and $(2)$ an institution cannot excessively decrease the capacity of division $j$ in response to increased demand in some divisions processed earlier.
@@
An aggregate choice rule is in the GS family if divisions’ choice rules satisfy the substitutes condition, size monotonicity, and quota monotonicity, and the transfer function is monotonic. Aggregate choice rules in the GS family may fail to satisfy the substitutes property and size monotonicity. In our Theorem 1, we show that each aggregate choice rule in the GS family has a completion that satisfies the substitutes property and size monotonicity (Hatfield and Kominers (2019)). An immediate corollary of Theorem 1 is that the generalized deferred acceptance mechanism (or cumulative offer mechanism) is the unique mechanism that is stable and strategy-proof.
@@
In practice, policymakers frequently implement choice rules characterized by intuitive definitions. The general classes of choice rules can be too complex to articulate effectively (Echenique (2007)). Choice rules in the GS family appear naturally in important real-world applications. For example, policy goals and affirmative action restrictions in India uniquely identify a particular choice function for reserve categories, the two-step meritorious horizontal choice rules (Sönmez and Yenmez (2022)). We show that an aggregate choice rule of an institution in which each division uses the two-step meritorious horizontal choice rule is in the GS family under any monotone transfer function (Proposition 1). In addition, various choice rules that are being implemented and a new one proposed by Hu and Zhang (2025) for admission to high schools in China are also in the GS family.
@@
Other prominent choice rules in the GS family include the reserve-based choice rules of Echenique and Yenmez (2015), the slot-specific priority choice rules of Kominers and Sönmez (2016), and the dynamic reserves choice rules of Aygün and Turhan (2020) and the multi-price choice rule of Greenberg, Pathak, and Sönmez (2024), among many others. GS choice rules offer considerable latitude to integrate complex diversity constraints. Moreover, the general framework we introduce is flexible to accommodate many other real-world matching applications, such as airline seat upgrades (Kominers and Sönmez (2016)), multidimensional reserves in Brazil (Aygün and Bó (2021)), high school admissions in China (Hu and Zhang (2025)), affirmative action implementations in Chilean school choice (Correa et al. (2021)), among others. Thus, the GS family provides a unified framework for market designers.
The paper is organized as follows. Section 2 discusses the related literature and contributions of our paper. Section 3 presents the model that formulates the general theory of sequential reserves for a single institution. Section 4 formulates the centralized matching problem with multiple institutions. Section 4 introduces the completion of a generalized sequential choice rule and thoroughly investigates its properties. The main result of the paper on stable and strategy-proof mechanisms through GS choice rules is presented in Section 4. Section 5 presents two important real-world applications of the GS choice rules: affirmative action in India and admissions to high schools in China. Finally, Section 6 concludes. All formal proofs are presented in the appendix.
2. Literature Review
This paper offers two important contributions to the literature. First, we develop a generalized theoretical framework for the implementation of soft reserve policies, within a single institution, where reserve categories’ choice rules transcend the selection based on simple baseline priority orderings. Second, we extend our analysis to centralized markets with multiple institutions, demonstrating that both stability and strategy-proofness are preserved when employing the comprehensive family of GS choice rules.
@@ 2. Literature Review
The paper closest to ours is Hatfield, Kominers, and Westkamp (2017) in which the authors study matching markets where institutions have multiple divisions that face capacity constraints similar to ours. The authors also study how hiring in one division affects the resources available for hiring in other divisions. Unlike our GS choice rules, the complexity of choice rules with flexible allotments they introduce could make them challenging to implement in practice. Their allotment functions require determining how many positions are allocated to each division given the available set of alternatives.
@@
This paper contributes to the existing literature on diversity constraints in matching market design. The idea of reserve was introduced in the seminal work by Hafalir, Yenmez, and Yildirim (2013) after Kojima (2012) pointed out the drawbacks of the quota policy. Ehlers, Hafalir, Yenmez, and Yildirim (2014) show that, with hard bounds, there might not exist assignments that satisfy standard fairness and non-wastefulness properties. To achieve fair and non-wasteful assignments, the authors propose to interpret constraints as soft bounds. There is now a rich market design literature on implementing diversity in matching markets. Some other examples are Kamada and Kojima (2015), Biró, Hassidim, Romm, Shorrer, and Sóvágó (2020), and Hassidim, Romm, and Shorrer (2021).
@@
This paper contributes to the growing body of literature on resource allocation problems in India. In Section 5.1, we examine resource allocation problems in India under complex affirmative action constraints as a key application. We show that employing the meritorious horizontal choice rule (Sönmez and Yenmez (2022)) within each division (or reserve category) in conjunction with monotone transfer functions produces choice rules within the GS family. Our formulation explicitly accounts for the de-reservation of vacant Other Backward Classes (OBC) positions, a consideration absent from Sönmez and Yenmez (2022). [^2] Aygün and Turhan (2022) examine the concurrent implementation of vertical reservations and OBC de-reservations, specifically focusing on the recently revamped admissions procedures for technical universities in India, as originally designed and implemented by Baswana, Chakrabarti, Chandran, Kanoria, and Patange (2019). Furthermore, our model allows individuals to express preferences over pairs of institution-position category, improving the applicability of the framework to real-world allocation scenarios. Echenique and Yenmez (2015) is the first paper to present Indian affirmative action as an example of controlled school choice. Subsequent work by Aygün and Turhan (2017) delves into the specific challenges in admissions to the Indian Institutes of Technologies (IITs). Aygün and Turhan (2020) study implementation of vertical reservations and de-reservations in the context of admissions to IITs where the authors introduced a family of dynamic reserves choice rules. The family of GS choice rules presented in the current paper nests these dynamic reserve choice rules. This generalization is of significance, as the dynamic reserve choice rules proposed by Aygün and Turhan (2020) were formulated on the basis of divisions’ q-responsive choice rules, which are inherently limited to incorporate horizontal reservations.
3. A General Theory of Sequential Reserves
Consider an institution $s$ that has a set of divisions $\mathcal{K}_{s}=\{1,...,K_{s}\}$. Each division $k\in\mathcal{K}_{s}$ has an associated choice rule $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ that specifies the contracts that division $k$ chooses given a set of alternatives and a capacity that may vary as a function of the number of vacant positions in the preceding divisions. We require that for a set of contracts $Y\subseteq\ X_{s}$ and a capacity of $\kappa\in\mathbb{Z}_{+}$, $\mid C_{s}^{k}(Y;\kappa)\mid\leq\kappa$.
Given a set contracts $Y\equiv Y^{1}\subseteq\ X$ , a capacity $\overline{q}_{s}$ for institution $s$, and a capacity $\overline{q}_{s}^{1}$ for division 1, we compute the set chosen $C_{s}(Y,\overline{q}_{s})$ sequentially in $K_{s}$ steps, where division $k$ chooses in step $k$, for $k=1,...,K_{s}$, as follows:
Step 1
$r_{1}=\overline{q}_{s}^{1}-\mid C_{s}^{1}(Y^{1};\overline{q}_{s}^{1})\mid$
be the number of vacant positions in division 1 and
Step k ($2\leq k\leq K_{s})$
be the number of unfilled positions in division $k$ and
The union of divisions’ choices is the institution’s chosen set. That is,
Transfer Policy
$q_{s}^{j}(\widetilde{r}_{1},...,\widetilde{r}_{j-1})\geq q_{s}^{j}(r_{1},...,r_{j-1}),\;\text{and}$
$\sum_{m=2}^{j}[q_{s}^{m}(\tilde{r}_{1},...,\widetilde{r}_{m-1})-q_{s}^{m}(r_{1},...,r_{m-1})]\leq\sum_{m=1}^{j-1}[\widetilde{r}_{m}-r_{m}].$
Conditions on Divisions’ Choice Rules
A choice rule $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ satisfies the substitutes property if, for all $q\in\mathbb{Z}_{+}$ and for all $x,y\in X$ and $Y\subseteq X$,
$\mid C_{s}^{k}(Y;q)\mid\leq\mid C_{s}^{k}(Y\cup\{x\};q)\mid.$
The substitutes condition and size monotonicity jointly imply the irrelevance of rejected contracts (IRC) condition (Aygün and Sönmez (2013)). A choice rule $C_{s}^{k}$ satisfies the IRC condition if, for any $Y\subseteq X$ and any $x\in X\setminus Y$,
Both the substitutes property and size monotonicity are standard conditions that impose restrictions on the choice function when the set of alternatives changes, while the capacity remains unchanged. The next property imposes conditions on the choice function when the capacity increases while the set of alternatives remains unchanged. This property is in the spirit of the expansion property of the choice rules defined in Chambers and Yenmez (2017).
Definition 1.
$C_{s}^{k}(Y,q)\subseteq C_{s}^{k}(Y,q+\Delta),$
and
$\mid C_{s}^{k}(Y,q+\Delta)\mid-\mid C_{s}^{k}(Y,q)\mid\leq\Delta.$
The QM property requires that (1) if there is an increase in capacity, the choice rule selects every contract it was choosing before the increase in capacity from any given set of contracts, and (2) if the capacity increases by $\Delta$, then the difference between the number of contracts chosen with the increase in capacity and the initial capacity cannot exceed $\Delta$. The first condition is the expansion property of Chambers and Yenmez (2017). The second condition, which is similar to the non-excessive reduction property of Westkamp (2013), restricts the magnitude of expansion when capacity is changed.
We are now ready to define a generalized sequential (GS) choice rule.
Definition 2.
An aggregate choice rule $C_{s}$ is an GS rule if it can be represented as a pair $\left(\left(C_{s}^{k}\right)_{k\in\mathcal{K}_{s}},q_{s}\right)$ of a collection of division choice rules and transfer policy, where each division’s choice rule satisfies the substitution condition, the size monotonicity and the quota monotonicity, and the transfer policy is monotone.
The family of GS choice rules is a parametrized family of aggregate choice rules derived from choice functions that satisfy both the substitutes condition and the size monotonicity. This parametrization stems from two primary sources. First, the transfer function provides one dimension of parametrization. Different monotonic transfer functions generate distinct aggregate choice rules within this family. Second, parametrization arises from the expansion of divisions’ choice rules while maintaining quota monotonicity. Within our theoretical framework, a division may be characterized by a sequence of choice rules, with each rule corresponding to a specific integer capacity. [^3] By defining the choice rule of division $k$ as $C_{s}^{k}:2^{X}\times\mathbb{Z}_{+}\longrightarrow 2^{X_{s}}$ we emphasize that capacity of division $k$ may vary. In a sense, this definition assigns a series of choice functions to division $k$, one for each given integer capacity. When the capacity of a substitutable and size monotonic choice rule is adjusted, the result is a new choice rule that also satisfies these properties. The quota monotonicity property establishes the precise relationship between these capacity-differentiated choice rules. It is important to note that a substitutable and size monotonic choice rule can be "expanded" to another such rule through multiple different ways, further enriching the parametric space of the GS family.
When each division’s choice rule is q-responsive, which implies the substitutes property, size monotonicity, and QM, the aggregate choice rule becomes a dynamic reserves choice rule introduced in Aygün and Turhan (2020). Therefore, the family of GS rules nests the family of dynamic reserves choice rules. An important example of a choice rule that satisfies the substitute property, size monotonicity, and QM but violates responsiveness is the meritorious horizontal choice rule introduced in Sönmez and Yenmez (2022). The GS family also nests the family of slot-specific priorities choice rules, introduced by Kominers and Sönmez (2016), because Aygün and Turhan (2020) shows that each slot-specific priorities choice rule can be represented as a dynamic reserves choice rule.
It is important to note that the aggregate choice rule in the GS family may fail to satisfy the substitute property and size monotonicity, even though the divisions’ choice rules satisfy them. However, a stable and strategy-proof mechanism design is still possible.
4. Stable and Strategy-proof Mechanism Design

### v3

# Matching with Generalized Sequential Choice Rules

@@ Abstract.
This paper studies a many-to-one matching between individuals and institutions where institutions comprise multiple divisions and face cross-divisional constraints. We introduce a parametrized family of choice rules, which we call generalized sequential (GSq), that encompasses many different choice rules encountered in practice and in market design literature. We show that the cumulative offer mechanism (COM) is the unique stable and strategy-proof mechanism in a matching problem where all institutions have a GSq choice rule. We present two real-world applications in which choice rules in the GSq family emerge naturally: affirmative action in India’s public higher educational institutions and government jobs and high school admissions in China.
@@ 1. Introduction
Many real-life allocation problems involve matching individuals with institutions that comprise multiple divisions (or reserve categories), such as firms, hospitals, or schools. In rationing healthcare resources (aziz2024efficient, pathak2024fair, and evren2025reserve) and the allocation of public goods such as seats in prestigious engineering schools (baswanaetal2019, aygun2022dereserve) and government jobs (sonmez2022affirmative) available resources are divided into different reserve categories, each with a subset of individuals eligible for that category. Divisions are endowed with their own choice rules.
In some applications, divisions’ choice rules are induced by a strict ranking of individuals and always select the q-best individuals, whenever available, where $q$ denotes the capacity. Such choice rules are called q-responsive. Similarly, in many rationing practices, resource units are assigned according to a baseline priority order and the eligibility of candidates in each reserve category.
However, in many other allocation problems, divisions’ choices are more complicated. To be concrete, consider a division with two positions that selects applicants based on merit scores and requires at least one position to be filled with a woman applicant. Suppose that the two top-scoring applicants are male, followed by a female applicant. This division selects the male applicant with the highest score, together with the female applicant, to meet the gender requirement. This kind of choice rule is not responsive because the two highest-scoring applicants were not chosen.
@@
Often, divisions are interconnected in that the number of positions available in one division depends on how many applicants other divisions hire. This creates cross-divisional constraints across the institution. For example, consider a high school that reserves slots for two middle schools in the same catchment area, while the remaining slots are open to all candidates, including students from other areas. Suppose that the school first fills the reserved slots for the two middle schools, followed by open slots. If there are not enough applicants from any of the middle schools, the remaining positions reserved for the middle school are allocated as open positions. [^1] This is an example of a soft reserve policy. There is also a hard reserve policy in which the remaining reserved positions are not made available to others; See hafalir2013effective and ehlers2014school for a detailed discussion on hard and sort reservation policies. As such, the capacity of the "open" positions depends on the number of vacancies in the reserved slots for the two middle schools. roth1999redesign were the first to address interdivisional constraints of this kind in matching theory. They introduced the notion of “reversion" in the context of the National Residency Matching Program (NRMP), which matches doctors with hospital programs. Reversion is a practice in which unfilled positions in one hospital program can be transferred to other programs, while respecting the hospital’s overall capacity limits. In their analysis, roth1999redesign identified reversion as a significant complication in the NRMP’s matching process.
@@
A significant body of market design literature examines the design and implementation of soft reserve policies in rationing and resource allocation problems. The specific method of soft reservation implementation has important distributional consequences. In the context of admissions to technical universities in India, aygun2022dereserve introduced a soft reserve policy that reverts unfilled slots reserved for particular groups to the open category, which is filled before reserve categories by updating the position categories’ capacities and re-running the choice rule from scratch. In a similar framework, aygun2025affirmative study forward transfer choice rules in which reserved slots that are not filled are allocated as an open category after the reserve categories are filled. hu2025verifiable presents various soft reserve policies currently in use for high school admissions in China. pathak2024fair introduced a smart reserve matching algorithm in the context of allocating life-saving medical resources.
This paper examines the choice problem faced by multi-divisional institutions that encounter cross-divisional constraints. We define a family of choice rules, which we call generalized sequential (GSq), that encompasses many different choice rules encountered in practice and in market design literature. We show that the generalized deferred acceptance (DA) is the unique stable and strategy-proof mechanism in a matching problem where all institutions have a GSq choice rule. [^2] In our setting, the term ”sequential” in GSq choice rules does not refer to a choice model with time component, as there is no time component in our framework. We consider static choice rules.
The GSq choice rules allow divisions to use substitutable and size-monotonic choice rules rather than constrain them to simpler $q$-responsive choice rules. In addition, these rules facilitate a more comprehensive soft reserve policies, as defined through transfer functions. Specifically, we present a parameterized family of choice rules constructed from substitutable choice rules that satisfy size monotonicity — two of the most prominent properties in the market design literature.
@@
The choice rules in the GSq family have the following structure: There is a predetermined order of precedence over divisions according to which the institution fills the positions of its divisions. Each division has an initial number of positions available that may be updated and is endowed with a series of choice rules, one for each (possible) integer capacity. The choice of a division depends on the set of available options (or contracts) and its capacity, both of which are determined by the choices made by the preceding divisions. The set of available options (or contracts) can be thought of as the set of remaining options from the choices of preceding divisions. The capacity of each division is a function of the number of unfilled positions in the preceding divisions and is given by an exogenously specified transfer function. The aggregate choice rule of an institution is then defined as the union of its divisions’ choice rules.
@@
The choice rule for each division satisfies the substitutes and the size monotonicity properties for each possible integer capacity. The substitutes property (kelso1982job and gul1999walrasian) and size monotonicity (alkan2003stable and hatfield2005matching) impose restrictions on the choice function when the set of available alternatives changes for a given capacity. A choice rule satisfies the substitute property if there are no two alternatives $x$ and $y$ that are complementary, in the sense that gaining access to $x$ makes $y$ desirable. A choice rule is size monotonic if it chooses weakly more alternatives whenever the set of available alternatives expands.
In our framework, a division’s capacity may be increased if there are vacancies in the preceding divisions. We, therefore, need to specify how a division can select from any given set of options under different capacities. When the capacity of a division increases, the division’s choice rules with initial and increased capacities can be regarded as two different choice rules. These choice rules are closely connected via the quota monotonicity property, which is similar to the notion of expansion defined in chambers2017choice. [^3] The quota monotonicity property requires consistency in choices when the capacity increases. It is a stronger condition than the expansion property of chambers2017choice. The choice rule with the increased capacity is an expansion of the choice rule with the initial capacity. The quota monotonicity property is satisfied if
@@
We formulate cross-division constraints via transfer functions. Given the order of precedence over the divisions and capacity of the first division, the capacity of a division is a function of the number of vacancies in the divisions that precede it. We require that the transfer function satisfies a mild condition, called monotonicity (westkamp2013analysis). A transfer function is monotone if:
  1. (1)
    whenever there are weakly more seats unfilled in every division preceding the $j^{th}$ division, weakly more slots should be available for the $j^{th}$ division, and
  2. (2)
    the capacity increase at each division cannot exceed the total number of vacant slots from the divisions preceding it. That is, no additional seats are created.
Condition (1) requires that the capacity function of each division is weakly increasing. Note that condition (2) is aggregate across divisions up to $j$, rather than a direct bound on the capacity change of a single division.
@@
An aggregate choice rule is in the GSq family if divisions’ choice rules satisfy the substitutes property, size monotonicity, and quota monotonicity, and the transfer function is monotonic. Aggregate choice rules in the GSq family may fail to satisfy the substitutes property and size monotonicity. In Theorem 1, we show that each aggregate choice rule in the GSq family satisfies a novel condition defined by hatfield2019hidden. An immediate corollary of Theorem 1 is obtained when combined with the results from hatfield2017restud and hatfield2019hidden: The generalized deferred acceptance (DA) mechanism is the unique stable mechanism that is strategy-proof.
@@
In practice, policymakers frequently implement choice rules characterized by intuitive definitions. The general classes of choice rules can be too complex to articulate effectively (echenique2007counting). Choice rules in the GSq family appear naturally in important real-world applications. For example, policy goals and affirmative action restrictions in India uniquely identify a particular choice function for reserve categories, the two-step meritorious horizontal choice rules (sonmez2022affirmative). We show that an aggregate choice rule of an institution, where each division employs the two-step meritorious horizontal choice rule, is in the GSq family under any monotone transfer function (Proposition 1). In addition, various choice rules that are being implemented and a new one proposed by hu2025verifiable for admission to high schools in China are also in the GSq family (Propositions 2-6).
@@
The GSq family nests other prominent choice rules introduced in the market design literature, including the reserve-based choice rules of echenique2015control, slot-specific priority choice rules of kominers2016matching, dynamic reserves choice rules of aygun2020dynamic and multi-price choice rule of greenberg2024redesigning, among many others. GSq choice rules offer considerable latitude to integrate complex diversity constraints. Moreover, the general framework we introduce is flexible to accommodate many other real-world matching applications, such as airline seat upgrades (kominers2016matching), multidimensional reserves in Brazil (aygun2021college), high school admissions in China (hu2025verifiable), affirmative action implementations in Chilean school choice (Correa et al. (2021)), among others. Thus, the GSq family provides a unified framework for market designers.
1.1. Literature Review
@@
hatfield2017stable also study matching markets, where institutions have multiple divisions that face capacity constraints similar to ours. The authors also study how hiring in one division affects the resources available for hiring in other divisions. Unlike our GSq choice rules, the complexity of choice rules with flexible allotments they introduce could make them challenging to implement in practice. Their allotment functions require determining how many positions are allocated to each division given the available set of alternatives.
@@
This paper contributes to the existing literature on diversity constraints in matching market design. The idea of reserve was introduced in the seminal work by hafalir2013effective after kojima2012school pointed out the drawbacks of the quota policy. ehlers2014school show that, with hard bounds, there might not exist assignments that satisfy standard fairness and non-wastefulness properties. To achieve fair and non-wasteful assignments, the authors propose to interpret constraints as soft bounds. There is now a rich market design literature on implementing diversity in matching markets. Some other examples are kamada2015efficient, biro2020need, and hassidim2021limits.
@@
This paper contributes to the growing body of literature on resource allocation problems in India. In Section 5.1, we examine resource allocation problems in India under complex affirmative action constraints as an application. We show that employing the meritorious horizontal choice rule (sonmez2022affirmative) within each division in conjunction with monotone transfer functions produces choice rules within the GSq family. Our formulation explicitly accounts for the de-reservation of vacant Other Backward Classes (OBC) positions, a consideration absent from sonmez2022affirmative.444aygun2022dereserve examine the concurrent implementation of vertical reservations and OBC de-reservations, specifically focusing on the recently revamped admissions procedures for technical universities (IITs) in India, as originally designed and implemented by baswanaetal2019. See also aygun2023priority for computationally simpler than and outcome equivalent to backward transfers choice rule. Furthermore, our model allows individuals to express preferences over pairs of institution-position category, improving the applicability of the framework to real-world allocation scenarios. echenique2015control is the first paper to present Indian affirmative action as an example of controlled school choice. Subsequent work by aygun2017large delves into the specific challenges in admissions to IITs.
aygun2020dynamic introduced a family of dynamic reserves choice rules, which were shown to nest the slot-specific priorities choice rules of kominers2016matching. The family of GSq choice rules nests dynamic reserve choice rules. This generalization is significant, as dynamic reserve choice rules were formulated using q-responsive divisional choice rules, which are inherently limited in incorporating horizontal reservations considered in sonmez2022affirmative.
The rest of the paper is organized as follows. Section 2 formulates the model of many-to-one matching with contracts. Section 3 introduces the family of GSq choice rules that are central to our analysis. Section 4 presents the main result of the paper on stable and strategy-proof mechanisms through GSq choice rules. Section 5 presents two real-world applications in which GSq choice rules are used or proposed: affirmative action in India and high school admissions in China. Finally, Section 6 concludes. All formal proofs are presented in the appendix.
2. Preliminaries